1. Suppose we have the following demand curve P=100-Q/2 calculate the elasticity of demand at point A(P=50) and B(P=80)
1
Expert's answer
2021-02-24T15:04:03-0500

The elasticity of demand is equal to:'

"E_d=\dfrac{d Q}{d P}\times \dfrac{P}{Q}"

The demand equation is P=100-Q/2. Therefore:

"dP=-\dfrac{1}{2}dQ\\[0.3cm]
\dfrac{dQ}{dP}=-2"

At P=50, the quantity demanded is:

"50=100-Q/2\\[0.3cm]
Q=100"

Therefore, the elasticity of demand is:

"E_d=-2\times\dfrac{50}{100}\\
E_d=\boxed{-1}"

At P=80, the quantity demanded is:

"80=100-Q/2\\[0.3cm]
Q=40"

Therefore, the price elasticity of demand is:

"E_d=-2\times \dfrac{80}{40}\\
E_d=\boxed{-4}"